The population (in millions) of a certain country can be approximated by the function: p(x) = 50 ∙ 1.05^x Where x is the number of years after 2000. Which of the following calculations will tell in what year the population can be expected to reach 200 million?
1
Expert's answer
2019-03-13T11:18:42-0400

Need to solve the equation

50 *1.05^x=200 →1.05^x =4 х= log1.05200/50 =ln4/ ln1.05 = 1.38629/ 0.04879=28.41340

In 2029, the population will reach 200 million